# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v2

@@ 4.2. Comparative Baseline
    • iTransformer (ICLR 2024) (Liu et al., 2023): A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
    • PatchTST (ICLR 2023) (Nie et al., 2023): This model divides the time series into patches and applies transformer-based operations on each patch independently. It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
    • Crossformer (ICLR 2023) (Zhang and Yan, 2023): This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
    • FEDformer (ICML 2022) (Zhou et al., 2022): A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
    • ETSformer (Woo et al., 2022): A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
    • Stationary (NeurIPS 2022) (Liu et al., 2022): A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
    • Autoformer (NeurIPS 2021) (Wu et al., 2021): A self-supervised transformer model that learns to forecast future values without relying on explicit labels. It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
@@
    • Rlinear (ICLR 2024) (Li et al., 2023): This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data. It is effective at capturing periodic patterns and works well with longer input sequences.
    • Dlinear (AAAI 2023) (Zeng et al., 2023): A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data. It serves as a simple yet effective baseline for comparison.
    • LightTS (Zhang et al., 2022): A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
@@
    • TiDE (ICLR 2023) (Das et al., 2023): A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
    • TimeNet (ICLR 2023) (Wu et al., 2023): A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
@@ 4.3. Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores , 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch (Paszke et al., 2019). For the sake of fairness, we used an open-source benchmark located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The open source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network. The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam (Kingma and Ba, 2015). We used Mean Absolute Error (MAE) and MSE in Equation 15 and Equation 16 as error indicators in the experiment where $y_{i}$ is actual value and $\hat{y}_{i}$ is predicted value.
(15) $MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
(16) $MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
4.4. Forecasting Result
@@
In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table 2. The results showcase the remarkable capabilities of the Time Evidence Fusion Network (TEFN) in long-term time series forecasting. TEFN consistently achieves state-of-the-art (SOTA) performance across various datasets and prediction horizons, highlighting its effectiveness and versatility in capturing complex temporal patterns. Notably, TEFN achieves this with a significantly smaller model size compared to models such as iTransformer (Liu et al., 2023) and PatchTST (Nie et al., 2023), demonstrating its efficiency and scalability.
@@
TEFN has significant advantages over other models in terms of prediction accuracy, as evidenced by the MSE and MAE metrics. Specifically, in terms of MSE, TEFN is $0.028$ smaller than iTransformer (Liu et al., 2023) and $0.025$ smaller than PatchTST (Nie et al., 2023) in the task of predicting a length of $L_{pred}=720$ on the ETTh1 dataset (Li et al., 2021). Similarly, in terms of MAE, TEFN once again showed a clear lead over iTransformer. TEFN was $0.008$ smaller than iTransformer (Liu et al., 2023) and $0.007$ smaller than PatchTST (Nie et al., 2023) in predicting a length of $L_{pred}=96$ on the Exchange dataset (Wu et al., 2021). These results highlight the effectiveness and robustness of TEFN in time series prediction, demonstrating its ability to provide more accurate predictions compared to contemporary methods.
(a) TEFN (Ours)
(b) PatchTST
(c) Crossformer
(d) TimesNet
(e) Dlinear
(f) Stationary
Figure 5. Visualization of Electricity-96 tasks
@@
We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models. We take the task of Electricity-96 (ecl, [n.d.]) as an example and selected typical models PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), TimesNet (Wu et al., 2023), Dlinear (Zeng et al., 2023) and Stationary (Liu et al., 2022) for visualization in Figure 5. From the visualization results, TEFN fits the details and trends of time series better.
4.5. Hyperparameter Sensitivity
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 6. It is worth noting that the order of magnitude of the MSE in Figure 6 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small. Full hyperparameter sensitivity experiment results are in the Table 3.
@@ 4.5. Hyperparameter Sensitivity
Figure 6. Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table 4. The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation. Random parameter selection within the traversal range of TEFN yields stable prediction results.
Table 4. The variance of model error for traversal learning rate and sample space
@@
(a) Channel BPA
(b) Time BPA
Figure 7. Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time. The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
4.6. Ablation Study
Table 5. The results of the ablation experiment
@@ 4.6. Ablation Study
TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$. In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table 5. Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error. Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (error of Electricty only varies $0.001$). It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting. We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table 5. In the aspect of all the results of ablation, TEFN can effectively improve the performance by combining the dimensions of time $T$ and channel $C$ in most cases. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation 17.
(17) $l_{C+T}\leqslant min(l_{C},l_{T})$
4.7. Efficiency Comparison
TEFN is a compact model compared to the parameters of the Transformer. In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison. In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample. Different models occupy varying amounts of GPU and memory. Therefore, we used the size of the saved binary model files as the model size. We visualized this with a bubble chart in Figure 8 . It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters. Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency. TEFN maintains almost Dlinear (Zeng et al., 2023) performance and almost PatchTST (Nie et al., 2023) accuracy, making it a performance error balanced model.
Figure 8. Model efficiency comparison. Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iterations.
It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices. As evidenced by the official code repository (https://github.com/ztxtech/Time-Evidence-Fusion-Network), TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM. This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer (Zhang and Yan, 2023) when deployed on portable computers.
4.8. Interpretability Analysis

### v3

@@ 4.2 Comparative Baseline
    • iTransformer (ICLR 2024) [48]: A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
    • PatchTST (ICLR 2023) [36]: This model divides the time series into patches and applies transformer-based operations on each patch independently. It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
    • Crossformer (ICLR 2023) [20]: This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
    • FEDformer (ICML 2022) [52]: A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
    • ETSformer [51]: A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
    • Stationary (NeurIPS 2022) [42]: A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
    • Autoformer (NeurIPS 2021) [45]: A self-supervised transformer model that learns to forecast future values without relying on explicit labels. It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
@@
    • Rlinear (ICLR 2024) [49]: This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data. It is effective at capturing periodic patterns and works well with longer input sequences.
    • Dlinear (AAAI 2023) [39]: A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data. It serves as a simple yet effective baseline for comparison.
    • LightTS [40]: A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
@@
    • TiDE (ICLR 2023) [50]: A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
    • TimesNet (ICLR 2023) [43]: A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
@@ 4.3 Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores, 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch [53]. For the sake of fairness, we used an open-source benchmark  [^1] The open-source benchmark Time-Series-Library located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The open-source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network.. The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam [54]. We used Mean Absolute Error (MAE) and MSE in Equation (15) and Equation (16) as error indicators in the experiment where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the predicted value.
$MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (15)
$MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (16)
4.4 Forecasting Result
@@
In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table II. The results showcase the remarkable capabilities of the Time Evidence Fusion Network (TEFN) in long-term time series forecasting. TEFN consistently achieves state-of-the-art (SOTA) performance across various datasets and prediction horizons, highlighting its effectiveness and versatility in capturing complex temporal patterns. Notably, TEFN achieves this with a significantly smaller model size compared to models such as iTransformer [48] and PatchTST [36], demonstrating its efficiency and scalability.
@@
TEFN has significant advantages over other models in terms of prediction accuracy, as evidenced by the MSE and MAE metrics. Specifically, in terms of MSE, TEFN is $0.028$ smaller than iTransformer [48] and $0.025$ smaller than PatchTST [36] in the task of predicting a length of $L_{pred}=720$ on the ETTh1 dataset [19]. Similarly, in terms of MAE, TEFN once again showed a clear lead over iTransformer. TEFN was $0.008$ smaller than iTransformer [48] and $0.007$ smaller than PatchTST [36] in predicting a length of $L_{pred}=96$ on the Exchange dataset [45]. These results highlight the effectiveness and robustness of TEFN in time series prediction, demonstrating its ability to provide more accurate predictions compared to contemporary methods.
(a) TEFN (Ours)-Electricity
(b) PatchTST-Electricity
(c) Crossformer-Electricity
(d) TimesNet-Electricity
(e) Dlinear-Electricity
(f) Stationary-Electricity
(g) TEFN (Ours)-ETTm1
(h) PatchTST-ETTm1
(i) Crossformer-ETTm1
(j) TimesNet-ETTm1
(k) Dlinear-ETTm1
(l) Stationary-ETTm1
Figure 5: Visualization of long-term forecasting tasks: Figure 5(a) - Figure 5(f) in Electricity dataset with $L_{pred}=96$ and Figure 5(g) - Figure 5(l) in ETTm1 dataset with $L_{pred}=96$
@@
We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models. We take the task of Electricity-96 [44] and ETTm1-96 as an example and selected typical models PatchTST [36], Crossformer [20], TimesNet [43], Dlinear [39] and Stationary [42] for visualization in Figure 5. From the visualization results, TEFN fits the details and trends of time series better.
4.5 Efficiency Comparison
TEFN is a compact model compared to the parameters of the Transformer. In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison. In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample. Different models occupy varying amounts of GPU and memory. Therefore, we used the size of the saved binary model files as the model size. We visualized this with a bubble chart in Figure 6 . It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters. Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency. TEFN maintains almost Dlinear [39] performance and almost PatchTST [36] accuracy, making it a performance error balanced model.
Figure 6: Model efficiency comparison. Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iteration. The models beyond the canvas have not been drawn.
It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices. As evidenced by the official code repository  [^2] https://github.com/ztxtech/Time-Evidence-Fusion-Network, TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM. This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer [20] when deployed on portable computers.
According to the process of BPA, the time complexity of BPA is $O(n*2^{|S|})$ where $n$ is input size and $|S|$ is sample space size. For the TEFN model, the specific time complexity is $O(C*L*2^{|S|})$ where $C$ is channel size, and $L$ is the length of time series. $C$ and $|S|$ are constants and satisfy $C,|S|<<L$. Hence, the final time complexity is $O(L)$. So TEFN can maintain positive performance in long-term forecasting tasks. Another factor affecting experimental efficiency is whether TEFN is linear or nonlinear. We chose to use a linear BPA module in the experimental section because of its high efficiency. Nonlinear modules also exist, and only the membership functions need to be replaced. BPA needs to correspond to different types of nonlinear time series and design membership functions. In the future, we will explore adaptive nonlinear functions. The fitting of nonlinear functions also brings higher computational costs, therefore, in the case of efficiency priority, the linear version of TEFN is the preferred choice.
4.6 Hyperparameter Sensitivity
@@
Figure 7: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 7. It is worth noting that the order of magnitude of the MSE in Figure 7 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small. In order to more intuitively reflect the hyperparameter sensitivity of TEFN, correlation analysis [55, 56]  [^3] The implementation of relevant calculations for discrete indicators such as Dataset comes from dython Package (https://shakedzy.xyz/dython/). is conducted on prediction length (Target), dataset, sample space (S), Learning rate (Lr), MAE and MSE in Figure 8. According to the correlation between different indicators, it can be found that TEFN is stable to the learning rate and sample space size which are almost uncorrelated with the two error indicators MSE and MAE.
Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table III. The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation. Random parameter selection within the traversal range of TEFN yields stable prediction results.
TABLE III: The variance of model error for traversal learning rate and sample space
@@
Figure 8: Heatmap of hyperparameter sensitivity correlation
4.7 Ablation Study
TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$. In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table IV. Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error. Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (the error of Electricity only varies $0.001$). It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting. We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table IV. In the aspect of all the results of ablation, TEFN can effectively improve performance by combining the dimensions of time $T$ and channel $C$ in most cases. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation (17).
$l_{C+T}\leqslant min(l_{C},l_{T})$ (17)
TABLE IV: The results of the ablation experiment
@@ 4.7 Ablation Study
(a) Channel BPA
(b) Time BPA
Figure 9: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time. The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
4.8 Robustness Experiments
TEFN is a highly robust neural network. To verify the robustness of TEFN in real prediction, a comparative experiment was conducted by adding noise to the dataset. The unit noise $\gamma$ follows the Gaussian distribution $N(0,1)$ with an expectation of 00 and a variance of $1$. Assume that the input time series is $x$, with time and channel dimensions. To adapt to the scale of the data, the transformation as Eq.(18) is applied to input data where $i$ is the index of channel dimension, $i$ is the index of time dimension, $x_{i}$ is a vector containing all time steps of channel $i$ and $std(\cdot)$ is Standard deviation function. So, the noisy data experiments are compared with the original data experiments.
$x^{\prime}_{ij}=x_{ij}+std(x_{i})*\gamma$ (18)
The experimental results of robustness testing are shown in the Table V. Continuing from the analysis, the experimental outcomes underscore the robust nature of TEFN, where the introduction of Gaussian noise to the dataset led to only marginal increases in prediction errors, with most variations being insignificant relative to the original errors. Notably, in long-term forecasting scenarios, a counterintuitive improvement in accuracy was observed in some instances, indicating that TEFN’s adaptability and learning capacity in complex, dynamic environments extend beyond mere resilience to noise, potentially leveraging subtle perturbations to enhance prediction performance.
TABLE V: The results of robustness experiments
Dataset Pred Original Noisy
MSE MAE MSE MAE
ETTm1 96 0.343 0.367 0.414 ($\uparrow$ 0.071) 0.411 ($\uparrow$ 0.044)
192 0.381 0.383 0.443 ($\uparrow$ 0.062) 0.429 ($\uparrow$ 0.046)
336 0.414 0.404 0.499 ($\uparrow$ 0.085) 0.458 ($\uparrow$ 0.054)
720 0.475 0.438 0.381 ($\downarrow$ 0.094) 0.398 ($\downarrow$ 0.040)
ETTm2 96 0.181 0.264 0.254 ($\uparrow$ 0.073) 0.312 ($\uparrow$ 0.048)
192 0.246 0.304 0.314 ($\uparrow$ 0.068) 0.350 ($\uparrow$ 0.046)
336 0.307 0.343 0.411 ($\uparrow$ 0.104) 0.403 ($\uparrow$ 0.060)
720 0.407 0.398 0.199 ($\downarrow$ 0.208) 0.281 ($\downarrow$ 0.117)
ETTh1 96 0.383 0.391 0.466 ($\uparrow$ 0.083) 0.448 ($\uparrow$ 0.057)
192 0.433 0.419 0.508 ($\uparrow$ 0.075) 0.468 ($\uparrow$ 0.049)
336 0.475 0.441 0.509 ($\uparrow$ 0.034) 0.484 ($\uparrow$ 0.043)
720 0.475 0.464 0.426 ($\downarrow$ 0.049) 0.427 ($\downarrow$ 0.037)
ETTh2 96 0.288 0.337 0.386 ($\uparrow$ 0.098) 0.401 ($\uparrow$ 0.064)
192 0.375 0.392 0.432 ($\uparrow$ 0.057) 0.441 ($\uparrow$ 0.049)
336 0.423 0.434 0.466 ($\uparrow$ 0.043) 0.467 ($\uparrow$ 0.033)
720 0.434 0.446 0.304 ($\downarrow$ 0.130) 0.352 ($\downarrow$ 0.094)
Electricity 96 0.197 0.273 0.270 ($\uparrow$ 0.073) 0.363 ($\uparrow$ 0.090)
192 0.197 0.276 0.281 ($\uparrow$ 0.084) 0.371 ($\uparrow$ 0.095)
336 0.212 0.292 0.322 ($\uparrow$ 0.110) 0.399 ($\uparrow$ 0.107)
720 0.253 0.325 0.270 ($\uparrow$ 0.017) 0.361 ($\uparrow$ 0.036)
Traffic 96 0.645 0.383 0.727 ($\uparrow$ 0.082) 0.479 ($\uparrow$ 0.096)
192 0.598 0.360 0.734 ($\uparrow$ 0.136) 0.481 ($\uparrow$ 0.121)
336 0.605 0.362 0.768 ($\uparrow$ 0.163) 0.491 ($\uparrow$ 0.129)
720 0.644 0.382 0.774 ($\uparrow$ 0.130) 0.494 ($\uparrow$ 0.112)
Weather 96 0.182 0.227 0.232 ($\uparrow$ 0.050) 0.269 ($\uparrow$ 0.042)
192 0.227 0.262 0.285 ($\uparrow$ 0.058) 0.305 ($\uparrow$ 0.043)
336 0.279 0.298 0.357 ($\uparrow$ 0.078) 0.350 ($\uparrow$ 0.052)
720 0.352 0.344 0.190 ($\downarrow$ 0.162) 0.237 ($\downarrow$ 0.107)
Exchange 96 0.082 0.198 0.188 ($\uparrow$ 0.106) 0.308 ($\uparrow$ 0.110)
192 0.176 0.297 0.340 ($\uparrow$ 0.164) 0.422 ($\uparrow$ 0.125)
336 0.328 0.412 0.963 ($\uparrow$ 0.635) 0.739 ($\uparrow$ 0.327)
720 0.861 0.700 0.092 ($\downarrow$ 0.769) 0.213 ($\downarrow$ 0.487)
4.9 Interpretability Analysis